OpenStudy (anonymous):

Find the linear approximation L (x) of the function F(x)=x^6+3x^2 at a=-1....L(x)=A+Bx where A=? and B=?

OpenStudy (anonymous):

L = f(a) + f'(a)(x-a)... L = x^6 + 3x^2 + (6x^5 + 6x)(x-a)... L = (-1)^6 + 3(-1)^2 + (6(-1)^5 +6(-1))(x-(-1))... L = 4 + (-12)(x+1)... L = 4 -12x -12... L = -8-12x...

OpenStudy (anonymous):

so how do u figure out what a and b equal

OpenStudy (bahrom7893):

Well wasn't it given: L = A + Bx quantish above solved: L = -8 - 12x L = A + Bx A = -8; B = -12